# Discrete Fourier Transform versus Discrete Chi-Square Method

### v2

@@ 4 Discussion
The solution for the non-linear DCM model is an ill-posed problem (Equations 1-5). We present a numerical solution that fulfils the $\mathrm{C}_{1}$, $\mathrm{C}_{2}$ and $\mathrm{C}_{3}$ conditions of a well-posed problem. The DCM model is just one special case of a non-linear model. Our technique can be applied to solve other non-linear models: Use the free parameters that make the model non-linear (Equation 9: $\bm{\beta}_{I}$) for solving the remaining other free parameters (Equation 2.1: $\bm{\beta}_{II}$).
@@ 4.1 Existence ( $\mathrm{C}_{1}$)
where the number of tested long and short search frequencies is $n_{\mathrm{L}}$ and $n_{\mathrm{S}}$, respectively. The number of bootstrap samples is $n_{\mathrm{B}}$. In other words, we solve this ill-posed problem by using brute numerical force. If the data contain only zero mean white noise, a linear least squares fit solution exists for every tested model.
@@ 4.3 Stability ( $\mathrm{C}_{3}$)
The artificial bootstrap data samples (Equation 22) represent “small changes in the input data”, while the bootstrap results for the model parameters represent “small chances in the solution” (Section 1: $\mathrm{C}_{3}$ condition formulation). We routinely check the stability of these solutions (e.g. Jetsu 2025, Figure S5). The unstable models, where the model parameter changes are large, are rejected (Section 2.1: $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{UM}}$}” models).
There are additional signatures of stability. The $z$ periodogram solution is unique for every tested $\bm{\beta}_{I}$ frequency combination. If these periodograms are continuous and their changes are not irregular (e.g., like in Figs. 1a-b), the DCM model solution is stable because it does not change by increasing the number of tested frequencies $n_{\mathrm{L}}$ and $n_{\mathrm{S}}$. Furthermore, the solutions for all seven simulated data samples converge when the $n$ and $\mathrm{SN}$} values increase (Tables 1 - 7). The different $n$ and $\mathrm{SN}$} combinations give the same stable DCM model solution.
We conclude that our numerical DCM model solution fulfils the $\mathrm{C}_{1}$, $\mathrm{C}_{2}$ and $\mathrm{C}_{3}$ conditions of the solution for a well-posed problem (Sections 4.1 - 4.3).
@@ 4.4 Performance
DCM does not suffer from the DFT limitations (Equations 26-29). The simulated Models 1-7 expose these DFT weaknesses, but in all these cases, DCM detects all periodic signals superimposed on an unknown trend (Sections 3.1 - 3.7). The correct signal(-s) and trend are detected. If there is enough $(n)$ very accurate $(\sigma_{i})$ data $(y_{i})$, DCM can predict the future or postdict the past by finding the correct DCM model for these data. We do admit that the $n$ and $\mathrm{SN}$} values of Models 2, 5 and 7 are extreme and unrealistic for most cases of real data. However, those model simulations are necessary for demonstrating that DCM can detect signals from short $\Delta T$ samples of high quality data. Those DCM signal detections depend only on the number of observations and their accuracy.
The above-mentioned Model 2, 5 and 7 simulations also reveal that the relative accuracy of amplitude and period estimates is lower than the relative accuracy of signal minimum and maximum epoch estimates (Tables 2, 5 and 7). This statistical effect is the same when DCM is applied to the data of any arbitrary phenomenon. This effect would, for example, explain why our solar cycle amplitude predictions are less accurate than our solar cycle minimum and maximum epoch predictions (Jetsu 2025). DCM detects the correct simulated period values, but DFT detects less accurate period values, which are not always correct. The leakage of DFT spectral power [20, 13] would explain why correct signal periods have not been detected earlier from the sunspot data.
We show that the DCM signal detections depend only on $n$ and $\mathrm{SN}$}, not on the time span $\Delta T$. This dimensional effect makes sense. If the data are inside an area $\Delta T\times\Delta y$, the ratio $y_{i}/\sigma_{i}$ reveals or hides the $g_{i}$ model details. DCM is an ideal forecasting method because the $\Delta T$ time span of data is irrelevant. This may motivate fast sampling of large high quality datasets and the use of numerical DCM parallel Python computation code in the analysis of those datasets.
4.5 Methodological contribution
What is our methodological contribution? DCM has already been published and DFT has been applied in innumerable studies. We would get the same results if DCM were compared to any other time series analysis method than DFT.
@@ 4.5 Methodological contribution
Practical time series analysis applications, like digital signal processing, require fast algorithms [20]. Our statistical numerical solution for the non-linear time series analysis model requires a lot of CPU. Yet, the CPU is beside the point if a well-posed numerical solution can be found. DCM proceeds through two stages. For every chosen alternative model, DCM performs a massive number of linear least squares fits (Equation 18) to find the unique initial free parameter values for the non-linear iteration (Equation 21). This first stage devours CPU. The second fast stage, the Fisher-test, then reveals the best model alternative (Equation 39).
DCM is an extremely simple time series analysis method that works in the time domain and uses real numbers. The more elaborate spectral estimating techniques, like DFT, work in the frequency domain and use complex numbers.
@@
We also want to draw attention to our idea of frequency space symmetry. This symmetry reduces the number of tested frequency combinations (Equation 11). For example, were this symmetry not exploited, the four signal $K_{1}=4$ search would give $4!=24$ same solutions in four-dimensional frequency space. The identification of the best frequency combination would be a daunting exercise, let alone the CPU consumed in testing all possible frequency combinations inside a tesseract (a four-dimensional cube). Our one-dimensional periodogram slices can "see inside” such multi-dimensional structures (Equation 20).
Generally, different time series analysis methods suffer from different limitations. We list some of those limitations below.
@@
DCM does not suffer from any of these limitations. This is our methodological contribution.
We hope that our results are considered to represent advances in the field of time series analysis. Our DCM can not fail if the Gauß-Markov theorem is valid.
@@
Consent for publication The author of this manuscript grants permission for the publication in Statistics and Computing.

### v3

@@ 4 Discussion
The solution for the non-linear DCM model is an ill-posed problem (Equations 1-5). We present a computational statistical solution that fulfils the $\mathrm{C}_{1}$, $\mathrm{C}_{2}$ and $\mathrm{C}_{3}$ conditions of a well-posed problem. The DCM model is just one special case of a non-linear model. Our technique can be applied to solve other non-linear models: Use the free parameters that make the model non-linear (Equation 9: $\bm{\beta}_{I}$) for solving the remaining other free parameters (Equation 2.1: $\bm{\beta}_{II}$).
@@ 4.1 Existence ( $\mathrm{C}_{1}$)
where the number of tested long and short search frequencies is $n_{\mathrm{L}}$ and $n_{\mathrm{S}}$, respectively. The number of bootstrap samples is $n_{\mathrm{B}}$. In other words, we solve this ill-posed problem by using brute computational force. If the data contain only zero mean white noise, the Gauß-Markov theorem ensures that a linear least squares fit solution exists for every tested frequency combination.
@@ 4.3 Stability ( $\mathrm{C}_{3}$)
The artificial bootstrap data sets (Equation 23) represent “small changes in the input data”, while the bootstrap results for the model parameters represent “small chances in the solution” (Section 1: $\mathrm{C}_{3}$ condition formulation). We routinely check the stability of these solutions (e.g. Jetsu 2025, Figure S5). The unstable models, where the model parameter changes are large, are rejected (Section 2.1: $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf\mathrm{UM}}$}” models).
There are additional signatures of stability. The $z$ periodogram solution is unique for every tested $\bm{\beta}_{I}$ frequency combination. If these periodograms are continuous and their changes are not irregular (e.g., like in Figs. 1a-b), the DCM model solution is stable because it does not change by increasing the number of tested frequencies $n_{\mathrm{L}}$ and $n_{\mathrm{S}}$. Furthermore, the solutions for all seven simulated data sets converge when the $n$ and $\mathrm{SN}$} values increase (Tables 1 - 7). The different $n$ and $\mathrm{SN}$} combinations give the same stable DCM model solution.
We conclude that our computational statistical DCM model solution fulfils the $\mathrm{C}_{1}$, $\mathrm{C}_{2}$ and $\mathrm{C}_{3}$ conditions of the solution for a well-posed problem (Sections 4.1 - 4.3).
@@ 4.4 Performance
DCM does not suffer from the DFT limitations (Equations 27-30). The simulated complex data sets of Models 1-7 expose these DFT weaknesses, but in all these cases, DCM detects all periodic signals superimposed on an unknown trend (Sections 3.1 - 3.7). The correct signal(-s) and trend are detected. If there is enough $(n)$ very accurate $(\sigma_{i})$ data $(y_{i})$, DCM can predict the future or postdict the past by finding the correct DCM model for these data. We do admit that the $n$ and $\mathrm{SN}$} values of Models 2, 5 and 7 are extreme and unrealistic for most cases of real data. However, those model simulations are necessary for demonstrating that DCM can detect signals from short $\Delta T$ samples of high quality data. Those DCM signal detections depend only on the number of observations and their accuracy.
Our DCM computes the $R$ (Equation 13) or $\chi^{2}$ (Equation 14) values for a massive number of linear models (Equation 46). The Gauß-Markov theorem ensures that the model having the lowest $R$ or $\chi^{2}$ is eventually always found, and the result for the non-linear iteration is unique (Equation 21). We could have simulated more complex data sets by increasing the values of $K_{1}$ (number of signals), $K_{2}$ (signal shape) and $K_{3}$ (polynomial order of trend). DCM would eventually detect those signals and the trend when the sampled data sets get larger $(n)$ and/or the signal to noise ratio becomes higher $({\mathrm{SN}})$. In this sense, DCM has no limit for the complexity of the data sets. In the real world, it may be impossible to eliminate noise $({\mathrm{SN}})$, but it is possible increase the sample size $(n)$. Since the time span $(\Delta T)$ is irrelevant, the best solution is fast sampling of accurate data sets. For example, many observers could measure the same phenomenon simultaneously. We recommend such fast sampling combined to DCM parallel computation Python code analysis.
We want to draw attention to three aspects of DCM computational performance.
  1. 1.
@@
    Practical time series analysis applications, like digital signal processing, require fast computational algorithms [22]. Our computational statistical solution for the non-linear DCM model requires a lot of CPU. Yet, the CPU used is beside the point if a well-posed computational solution can be found to any challenging scientific non-linear problem. DCM proceeds through two stages. For every chosen alternative model, DCM performs a massive number of linear least squares fits (Equation 18) to find the unique initial free parameter values for the non-linear iteration (Equation 21). This first stage devours CPU. The second fast stage, the Fisher-test, then reveals the best model alternative (Equation 40).
  2. 2.
@@
    Our frequency space symmetry idea dramatically reduces CPU (Equation 11). For example, were this symmetry not exploited, the four signal $K_{1}=4$ search would give $4!=24$ exactly the same solutions from four-dimensional tested grid in frequency space. This would cause unnecessary use of CPU in the preliminary testing all possible frequency combinations inside a tesseract (a four-dimensional cube). After one necessary iteration of Equation 21, and 23 unnecessary ones, the identification of the best frequency combination would be a daunting exercise.
  3. 3.
    The DCM signal detections depend only on the sample size $n$ and the signal to noise ratio $\mathrm{SN}$}, but not on the time span $\Delta T$. This dimensional effect makes sense. If the $t_{i}$ and $y_{i}$ data are inside a rectangle $\Delta T\times\Delta y=(t_{n}-t_{1})\times[{\mathrm{max}}(y_{i})-{\mathrm{min}}(y_{i})]$, then the $y_{i}/\sigma_{i}$ ratio reveals or hides the $g_{i}$ model details. Since the time span of data $\Delta T$ is irrelevant, DCM is an ideal forecasting method. We have been deluded into unnecessary waiting for the repetition of the signal(-s). DCM can perceive the future and the past much “earlier” (from shorter $\Delta T$) than previously thought, like in Equations 27 and 29. This should motivate fast sampling of large high quality data sets and their analysis using the DCM parallel Python code. The invention of computers has made it possible to use this computational statistical method, which is probably the simplest possible one.
The analytical solution for the non-linear DCM model is ill-posed (Equation 1). Regardless of the data set complexity, DCM can find a well-posed computational statistical solution for this problem (Equation 21). So far, we have compared DCM only to DFT. Both methods are parametric and operate in the frequency-domain. There are innumerable similar parametric time series analysis methods [e.g. 22, 15]. Each one of these other time series analysis methods suffers from their own particular application limitations. We list some of those application limitations below.
@@
DCM does not suffer from any of these application limitations. This simple computational statistical method can fail if, and only if, the Gauß-Markov theorem is not valid. That is impossible.
@@
Consent for publication The author of this manuscript grants permission for the publication in Computational Statistics.